THE EFFECT OF REDUCED MASS ON CROSSED BEAM ROVIBRATIONAL ENERGY TRANSFER FROM S_{1} GLYOXAL (0^{0}, K′ = 0)

\underline{S.M. CLEGG}, C.S. PARMENTER, Department of Chemistry, Indiana University, Bloomington, IN 47405; S.-P. LU, Innovative Lasers Corp., 3280 East Hemisphere Loop #120, Tucson, AZ 85706.

Energy transfer from S_{1} glyoxal (CHO-CHO) in collision with additional target gases is now added to a list of collision partners (H_{2}, D_{2}, He, Kr, and C_{6}H_{12}) previously presented. A laser is used to pump glyoxal to the 0^{0}, 0 ≤ J′ ≤ 10 K′ = 0 level of the S_{1} (^{1}A_{u}) state. Rotational excitation with resolution of K states can be observed in competition with rovibrational energy transfer to the lowest frequency mode, ν_{7} (also seen with K state resolution). Relative cross sections are obtained from computer simulations of the observed inelastic scattering fluorescence spectrum. A more complete picture of the relationship between rovibrational energy transfer and the reduced mass of the collision partner will be presented. These results can also be used to further establish benchmarks for the 3-dimensional quantal scattering model that has so successfully been used with past glyoxal energy transfer studies^{a}.

aG.-J. Kroes, R.P.H. Rettschnick, and D.C. Clary, Chem. Phys. 148, 359 (1990).